3.20 Downloading Apps for Your Smartphone A
random sample of n = 461 smartphone users in
the US in January 2015 found that 355 of them
have downloaded an app.10
(a) Give notation for the parameter of interest, and
define the parameter in this context.
(b) Give notation for the quantity that gives the best
estimate and give its value.
(c) What would we have to do to calculate the
parameter exactly?
Answer #1

Part a.

p = parameter is all smartphone users in us,

Part b.

Proportion = phat=355/461

Part c.

The population is all smartphone users in the US. The population parameter of interest is p,the proportion of smartphone users who have downloaded an app.10, For this sample phat= 355/461 =0.77. Unless we have additional information, the best point estimate of the population parameter p is phat=0.77.

Hope this will be helpful. Thanks and God Bless You:) If any query in the solution, please post in comments